1. What volume of a 6.0 mol/L sulfuric acid solution would be required to make 500 mL of 0.24 mol/L sulfuric acid?
1
Expert's answer
2021-04-29T07:18:30-0400

Molar Mass of sulfuric acid = 98.079

6 × 98.079 = 588.474

0.24 × 0.5 = 6.0 × X

0.12= 6X

X =0.12/6

= 0.02L

= 20ml.